# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 5.1. Blowing up analysis
First, we recall the maximum principle for the $L^{\infty}$ of ${\bf{d}}$ from the transported Ginzburg–Landau heat flow:
@@
(5.1) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{d}}}+{{\bf{u}}}\cdot\nabla{{\bf{d}}}=\Delta{{\bf{d}}}-f({{\bf{d}}}),\\ \nabla\cdot{\bf{u}}=0,\\ {{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{array}\right.$
Lemma 5.1 ([39]).
@@
For $T>0$, assume that ${{\bf{u}}}\in L^{2}(0,T;V)$ and ${{\bf{d}}}_{0}\in H^{1}(\mathbb{T};\mathbb{R}^{3})$ with $|{{\bf{d}}}_{0}(x)|\leq 1$ a.e. $x\in\mathbb{T}^{3}$. Suppose ${{\bf{d}}}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3},\mathbb{R}^{3}))$ is a weak solution to (5.1). Then $|{{\bf{d}}}(x,t)|\leq 1$ for a.e. $(x,t)\in Q_{T}$.
@@
$\Theta_{z_{0}}(\tau_{0}r)=(\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})+\left((\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}|\pi|^{\frac{3}{2}}\right)^{2}\leq\frac{1}{2}\max\{\Theta_{z_{0}}(r),C_{0}r^{3}\}.$
@@
(5.2) $\left\{\begin{array}[]{l}|{{\bf{z}}}|\leq M_{0},,\quad|{{\bf{d}}}|\leq M_{0}\text{ in }\mathcal{Q}_{r_{i}}(z_{i}),\\ \Theta_{z_{i}}(r_{i})=\varepsilon_{i}^{3},\\ \Theta_{z_{i}}(\tau r_{i})>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right.$
@@
It is straightforward to verify that $({{\bf{z}}}^{i},{{\bf{v}}}^{i},{{\bf{d}}}^{i},{\pi}^{i})$ satisfies the re-scaled equation
(5.3) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla({{\bf{z}}}^{i}+{{\bf{v}}}^{i})+\nabla\pi^{i}\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad=\Delta{\bf{v}}^{i}-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}_{i}\odot\nabla{{\bf{d}}}^{i}-\frac{1}{2}|\nabla{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-r_{i}^{2}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \nabla\cdot{{\bf{v}}}^{i}=0,\\ \partial_{t}{{\bf{d}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla{{\bf{d}}}^{i}=\Delta{{\bf{d}}}^{i}-r^{2}_{i}f({{\bf{d}}}^{i}).\end{array}\right.$
And (5.2) becomes
(5.4) $\left\{\begin{array}[]{l}\displaystyle\int_{\mathcal{Q}_{1}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\int_{\mathcal{Q}_{1}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{2}=\varepsilon_{i}^{3},\\ \displaystyle\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{3}>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right.$
Then we define the blowing-up sequence
@@
(5.5) $\left\{\begin{array}[]{l}\partial_{t}\widehat{{\bf{v}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot\nabla(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})+\nabla\widehat{\pi}^{i}=\Delta\widehat{{\bf{v}}}^{i}\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathop{\rm div}\nolimits\left(\varepsilon_{i}\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{\varepsilon_{i}}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \mathop{\rm div}\nolimits\widehat{{\bf{v}}}^{i}=0,\\ \partial_{t}\widehat{{\bf{d}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{\bf{v}}^{i})\cdot\nabla\widehat{{\bf{d}}}^{i}=\Delta\widehat{{\bf{d}}}^{i}-\frac{r_{i}^{2}}{\varepsilon_{i}}f({\bf{d}}^{i}).\end{array}\right.$
@@
(5.6) $\left\{\begin{array}[]{l}\int_{\mathcal{Q}_{1}(0)}\widehat{{\bf{d}}}^{i}=0,\\ \int_{\mathcal{Q}_{1}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\int_{\mathcal{Q}_{1}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}=1,\\ \tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}>\frac{1}{2}\max\{1,C_{i}\left(\frac{r_{i}}{\varepsilon_{i}}\right)^{3}\}.\end{array}\right.$
From (5.6), by weak compactness, we can show there exists
@@
such that, after passing to a subsequence,
$(\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i},\widehat{\pi}^{i})\rightharpoonup(\widehat{{\bf{v}}},\widehat{{\bf{d}}},\widehat{\pi})\text{ in }L^{3}(\mathcal{Q}_{1}(0))\times L_{t}^{3}W_{x}^{1,3}(\mathcal{Q}_{1}(0))\times L^{\frac{3}{2}}(\mathcal{Q}_{1}(0)).$
It follows from (5.6) and the lower semicontinuity that
@@
$\sup_{i}\left(\|\widehat{{\bf{v}}}^{i}\|_{L_{t}^{\infty}L_{x}^{2}\cap L_{t}^{2}H_{x}^{1}(\mathcal{Q}_{1}(0))}+\|\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}H_{x}^{1}\cap L_{t}^{2}H_{x}^{2}(\mathcal{Q}_{1}(0))}\right)<\infty.$
@@
Define $\varphi_{i}(x,t)=\varphi\left(\frac{x-x_{i}}{r_{i}},\frac{t-t_{i}}{r_{i}^{2}}\right)$, $\forall(x,t)\in\mathbb{T}^{3}\times(0,\infty).$ Replacing $\varphi$ by $\varphi_{i}^{2}$ in the local energy inequality (4.1) we can show that
$\displaystyle\sup_{t_{i}-\frac{r_{i}^{2}}{4}\leq t\leq t_{i}}\int_{B_{r_{i}}(x_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)\varphi_{i}^{2}{\,\mathrm{d}x}+\int_{\mathcal{Q}_{r_{i}/2}(z_{i})}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2}\right)\varphi_{i}^{2}$
$\displaystyle\leq C\int_{\mathcal{Q}_{r_{i}}(z_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)|\partial_{t}\left(\varphi_{i}^{2}\right)|$
@@
By rescaling and Young’s inequality, we can show that
$\displaystyle\sup_{-\frac{1}{4}\leq t\leq 0}\int_{B_{\frac{1}{2}(0)}}(|\widehat{{\bf{v}}}^{i}|^{2}+|\nabla\widehat{{\bf{d}}}_{i}^{2}|^{2})+\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|\nabla\widehat{{\bf{v}}}^{i}|+|\nabla^{2}\widehat{{\bf{d}}}^{i}|^{2})\leq C,\qquad\forall i.$
Hence the claim holds. From Aubin–Lions’s lemma, we infer that
$(\widehat{{\bf{v}}}^{i},\widehat{{\bf{d}}}^{i})\to(\widehat{{\bf{v}}},\widehat{{\bf{d}}})\text{ strongly in }L^{3}(\mathcal{Q}_{\frac{1}{2}}(0))\times L_{t}^{3}W_{x}^{1,3}(\mathcal{Q}_{\frac{1}{2}}(0)).$
We observe that $(\widehat{{\bf{v}}},\widehat{{\bf{d}}},\widehat{\pi})$ solves the liner system
@@
(5.7) $\left\{\begin{array}[]{l}\partial_{t}\widehat{{\bf{v}}}+\nabla\widehat{\pi}=\Delta\widehat{{\bf{v}}},\\ \nabla\cdot\widehat{{\bf{v}}}=0,\\ \partial_{t}\widehat{{\bf{d}}}=\Delta\widehat{{\bf{d}}}.\end{array}\right.$
By the standard interior regularity estimate we have that $(\widehat{\textbf{u}},\widehat{\textbf{d}})\in C^{\infty}(\mathcal{Q}_{\frac{1}{4}}(0))$, $\widehat{\pi}\in L^{\infty}([-\frac{1}{16},0],C^{\infty}(B_{\frac{1}{4}}(0)))$ with the following estimate:
@@
(5.8) $\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{\textbf{u}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}$
@@
$\leq C\tau^{3}\left[\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|\widehat{\textbf{u}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\int_{\mathcal{Q}_{\frac{1}{2}}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}\right]$
@@
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{\textbf{u}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3})=\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{{\bf{v}}}|^{3}+|\nabla\widehat{{\bf{d}}}|^{3})+\tau^{-2}o(i),$
where $o(i)\to 0$ as $i\to\infty$. It remains to estimate the pressure term $\widehat{\pi}^{i}$. Taking the divergence in (5.5)${}_{1}$, we see that
(5.9) $-\Delta\widehat{\pi}^{i}=\varepsilon_{i}\mathop{\rm div}\nolimits^{2}\left[\widehat{\textbf{u}}^{i}\otimes\widehat{\textbf{u}}^{i}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right].$

### v2

@@ 5.1 Blowing up analysis
First, we recall the maximum principle [69, Lemma 2.2] for the $L^{\infty}$ of ${\bf{d}}$ from the transported Ginzburg–Landau heat flow :
@@
$\left\{\begin{array}[]{l}\partial_{t}{{\bf{d}}}+{{\bf{u}}}\cdot\nabla{{\bf{d}}}=\Delta{{\bf{d}}}-f({{\bf{d}}}),\\ \nabla\cdot{\bf{u}}=0,\\ {{\bf{d}}}(0)={{\bf{d}}}_{0}.\end{array}\right$. (45)
Lemma 4.
@@
For $T>0$, assume that ${{\bf{u}}}\in L^{2}(0,T;V)$ and ${{\bf{d}}}_{0}\in H^{1}(\mathbb{T};\mathbb{R}^{3})$ with $|{{\bf{d}}}_{0}(x)|\leq 1$ a.e. $x\in\mathbb{T}^{3}$. Suppose ${{\bf{d}}}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3},\mathbb{R}^{3}))$ is a weak solution to (45). Then $|{{\bf{d}}}(x,t)|\leq 1$ for a.e. $(x,t)\in Q_{T}$.
@@
$\displaystyle\Theta_{z_{0}}(\tau_{0}r)$ $\displaystyle=(\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}(|{{\bf{v}}}|^{3}+|\nabla{{\bf{d}}}|^{3})+\left((\tau_{0}r)^{-2}\int_{\mathcal{Q}_{\tau_{0}r}(z_{0})}|\pi|^{\frac{3}{2}}\right)^{2}$
$\displaystyle\leq\frac{1}{2}\max\{\Theta_{z_{0}}(r),C_{0}r^{3}\}.$
@@
$\left\{\begin{array}[]{l}|{{\bf{z}}}|\leq M_{0},,\quad|{{\bf{d}}}|\leq M_{0}\text{ in }\mathcal{Q}_{r_{i}}(z_{i}),\\ \Theta_{z_{i}}(r_{i})=\varepsilon_{i}^{3},\\ \Theta_{z_{i}}(\tau r_{i})>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right$. (46)
@@
It is straightforward to verify that $({{\bf{z}}}^{i},{{\bf{v}}}^{i},{{\bf{d}}}^{i},{\pi}^{i})$ satisfies the rescaled equation
$\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla({{\bf{z}}}^{i}+{{\bf{v}}}^{i})+\nabla\pi^{i}\\ \qquad\qquad\qquad=\Delta{\bf{v}}^{i}-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}_{i}\odot\nabla{{\bf{d}}}^{i}-\frac{1}{2}|\nabla{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-r_{i}^{2}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \nabla\cdot{{\bf{v}}}^{i}=0,\\ \partial_{t}{{\bf{d}}}^{i}+({{\bf{z}}}^{i}+{{\bf{v}}}^{i})\cdot\nabla{{\bf{d}}}^{i}=\Delta{{\bf{d}}}^{i}-r^{2}_{i}f({{\bf{d}}}^{i}).\end{array}\right$. (47)
And (46) becomes
$\left\{\begin{array}[]{l}\displaystyle\int_{\mathcal{Q}_{1}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\int_{\mathcal{Q}_{1}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{2}=\varepsilon_{i}^{3},\\ \displaystyle\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\bf{v}}}^{i}|^{3}+|\nabla{{\bf{d}}}^{i}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\pi^{i}|^{\frac{3}{2}}\right)^{3}>\frac{1}{2}\max\{\varepsilon_{i}^{3},C_{i}r_{i}^{3}\}.\end{array}\right$. (48)
Then we define the blow-up sequence
@@
$\left\{\begin{array}[]{l}\partial_{t}\widehat{{\bf{v}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot\nabla(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})+\nabla\widehat{\pi}^{i}\\ \quad\quad\quad=\Delta\widehat{{\bf{v}}}^{i}-\mathop{\rm div}\nolimits\left(\varepsilon_{i}\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{\varepsilon_{i}}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}}F({{\bf{d}}}^{i}){\rm I}_{3}\right),\\ \mathop{\rm div}\nolimits\widehat{{\bf{v}}}^{i}=0,\\ \partial_{t}\widehat{{\bf{d}}}^{i}+\varepsilon_{i}(\widehat{{\bf{z}}}^{i}+\widehat{\bf{v}}^{i})\cdot\nabla\widehat{{\bf{d}}}^{i}=\Delta\widehat{{\bf{d}}}^{i}-\frac{r_{i}^{2}}{\varepsilon_{i}}f({\bf{d}}^{i}).\end{array}\right$. (49)
@@
$\left\{\begin{array}[]{l}\int_{\mathcal{Q}_{1}(0)}\widehat{{\bf{d}}}^{i}=0,\\ \int_{\mathcal{Q}_{1}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\int_{\mathcal{Q}_{1}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}=1,\\ \tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\right)^{2}\\ \quad\quad>{\frac{1}{2}\max\left\{1,C_{i}\left(\frac{r_{i}}{\varepsilon_{i}}\right)^{3}\right\}.}\end{array}\right$. (50)
From (50), by weak compactness, we can show there exists
@@
such that, after passing to the limit upon a subsequence, we have
[…]
It follows from (50) and the lower semicontinuity that
@@
$\sup_{i}\left(\|\widehat{{\bf{v}}}^{i}\|_{L_{t}^{\infty}L_{x}^{2}\cap L_{t}^{2}H_{x}^{1}({\mathcal{Q}_{1/2}(0)})}+\|\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}H_{x}^{1}\cap L_{t}^{2}H_{x}^{2}({\mathcal{Q}_{1/2}(0)})}\right)<\infty.$
@@
Define $\varphi_{i}(x,t)=\varphi\left(\frac{x-x_{i}}{r_{i}},\frac{t-t_{i}}{r_{i}^{2}}\right)$, $\forall(x,t)\in\mathbb{T}^{3}\times(0,\infty).$ Replacing $\varphi$ by $\varphi_{i}^{2}$ in the local energy inequality (81) we can show that
$\displaystyle\sup_{t_{i}-\frac{r_{i}^{2}}{4}\leq t\leq t_{i}}\int_{B_{r_{i}}(x_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)\varphi_{i}^{2}{\,\mathrm{d}x}$
$\displaystyle+\int_{\mathcal{Q}_{r_{i}/2}(z_{i})}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2}\right)\varphi_{i}^{2}$
$\displaystyle\leq$ $\displaystyle C\int_{\mathcal{Q}_{r_{i}}(z_{i})}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)|\partial_{t}\left(\varphi_{i}^{2}\right)|$
@@
By rescaling and Young’s inequality, we can show that for every $i$,
$\displaystyle\sup_{-\frac{1}{4}\leq t\leq 0}\int_{B_{\frac{1}{2}(0)}}(|\widehat{{\bf{v}}}^{i}|^{2}+|\nabla\widehat{{\bf{d}}}^{i}|^{2})+\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|\nabla\widehat{{\bf{v}}}^{i}|+|\nabla^{2}\widehat{{\bf{d}}}^{i}|^{2})\leq C.$ (51)
Hence the claim holds. Now we intend to apply the Aubin–Lions lemma; up to now what we still need is to obtain estimates on $\partial_{t}\widehat{{\bf{v}}}^{i}$ and $\partial_{t}\widehat{{\bf{d}}}^{i}$. To achieve this goal, firstly, we observe that from the equations for $\widehat{{\bf{v}}}^{i}$ in (49)1,2, we can show that, for any test functions $\zeta\in C_{c,\rm{div}}^{\infty}(B_{\frac{1}{2}(0)};\mathbb{R}^{3})$ we have
@@
$\left|\int_{Q_{\frac{1}{2}}(0)}\partial_{t}\widehat{{\bf{v}}}^{i}\cdot\zeta\right|$ $\leq\varepsilon_{i}\left|\int_{Q_{\frac{1}{2}}(0)}(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot((\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\cdot\nabla\zeta)\right|+\left|\int_{Q_{\frac{1}{2}}(0)}\nabla\widehat{{\bf{v}}}^{i}:\nabla\zeta\right|$ (52)
@@
$+\left|\int_{Q_{\frac{1}{2}(0)}}\varepsilon_{i}\left(\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{\varepsilon_{i}}{2}|\nabla\widehat{{\bf{d}}}^{i}{\rm I_{3}}-\frac{r_{i}^{2}}{\varepsilon_{i}}F({{\bf{d}}}^{i}){\rm I}_{3}\right):\nabla\zeta\right|$
$\leq C\|\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\|_{L^{\infty}_{t}L_{x}^{2}(Q_{\frac{1}{2}}(0))}^{\frac{1}{2}}\|\ \widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\|^{\frac{3}{2}}_{L_{t}^{2}L_{x}^{6}(Q_{\frac{1}{2}}(0))}\|\nabla\zeta\|_{L_{t}^{4}{L_{x}^{2}}(Q_{\frac{1}{2}}(0))}$
$+C\|\widehat{{\bf{v}}}^{i}\|_{L_{t}^{2}H_{x}^{1}(Q_{\frac{1}{2}}(0))}\|\nabla\zeta\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$+C\|\nabla\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}L_{x}^{2}(Q_{\frac{1}{2}}(0))}^{\frac{1}{2}}\|\nabla\widehat{{\bf{d}}}^{i}\|_{L_{t}^{2}L_{x}^{6}(Q_{\frac{1}{2}}(0))}^{\frac{3}{2}}\|\nabla\zeta\|_{L_{t}^{4}{L_{x}^{2}}(Q_{\frac{1}{2}}(0))}$
$+C\|F({{\bf{d}}}^{i})\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\nabla\zeta\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$\leq C\|\zeta\|_{L_{t}^{4}{\bf V}(Q_{\frac{1}{2}}(0))},$
where we have used the uniform estimates already obtained in (51), together with the uniform $L^{\infty}$ bounds for ${{\bf{d}}}$ and ${{\bf{z}}}$ provided in (46). Then from (52) we obtain the following uniform estimate:
$\|\partial_{t}\widehat{{\bf{v}}}^{i}\|_{L_{t}^{\frac{4}{3}}{\bf V}^{\prime}(Q_{\frac{1}{2}}(0))}\leq C,\quad\forall i=1,2,,\cdots.$ (53)
Now for the equations for $\widehat{{\bf{d}}}^{i}$ in (49)3, similarly, we can conclude that for any test function $\phi\in C_{0}^{\infty}(B_{\frac{1}{2}}(0))$ we have
@@
$\left|\int_{Q_{\frac{1}{2}}(0)}\partial_{t}\widehat{{\bf{d}}}^{i}\cdot\phi\right|$ $\leq\varepsilon_{i}\left|\int_{Q_{\frac{1}{2}}(0)}((\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i})\otimes\widehat{{\bf{d}}}^{i}):\nabla\phi\right|+\left|\int_{Q_{\frac{1}{2}}(0)}\nabla\widehat{{\bf{d}}}^{i}:\nabla\phi\right|$
$+\frac{r_{i}^{2}}{\varepsilon_{i}}\left|\int_{Q_{\frac{1}{2}}(0)}f({{\bf{d}}}^{i})\cdot\phi\right|$
$\leq C\|\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\widehat{{\bf{d}}}^{i}\|_{L_{t}^{\infty}L_{x}^{\infty}(Q_{\frac{1}{2}}(0))}\|\nabla\phi\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$+C\|\nabla\widehat{{\bf{d}}}^{i}\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\nabla\phi\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$+C\|f({{\bf{d}}}^{i})\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}\|\phi\|_{L_{t}^{2}L_{x}^{2}(Q_{\frac{1}{2}}(0))}$
$\leq C\|\phi\|_{L_{t}^{2}H_{x}^{1}(Q_{\frac{1}{2}}(0))}.$
This implies that
$\|\partial_{t}\widehat{{\bf{d}}}^{i}\|_{L_{t}^{2}H_{x}^{-1}(Q_{\frac{1}{2}}(0))}\leq C,\quad\forall i=1,2,\cdots.$ (54)
With the estimates on $\partial_{t}\widehat{{\bf{v}}}^{i}$ and $\partial_{t}\widehat{{\bf{d}}}^{i}$ in hand, now we are ready to apply the Aubin–Lions lemma. We find that
[…]
By the standard interior regularity estimate we can conclude that $({{\widehat{{\bf{v}}}}},\widehat{\textbf{d}})\in C^{\infty}(\mathcal{Q}_{\frac{1}{4}}(0))$ and $\widehat{\pi}\in L^{\infty}([-\frac{1}{16},0],C^{\infty}(B_{\frac{1}{4}}(0)))$ with the following estimate:
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\widehat{{\bf{v}}}}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}$ (56)
$\leq C\tau^{3}\left[\int_{\mathcal{Q}_{\frac{1}{2}}(0)}(|{{\widehat{{\bf{v}}}}}|^{3}+|\nabla\widehat{\textbf{d}}|^{3})+\left(\int_{\mathcal{Q}_{\frac{1}{2}}(0)}|\widehat{\pi}|^{\frac{3}{2}}\right)^{2}\right]$
@@
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|{{\widehat{{\bf{v}}}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3})=\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}(|\widehat{{\bf{v}}}|^{3}+|\nabla\widehat{{\bf{d}}}|^{3})+\tau^{-2}o(i),$
where $o(i)\to 0$ as $i\to\infty$. It remains to estimate the pressure term $\widehat{\pi}^{i}$. Taking the divergence in (49)1, we see that
$-\Delta\widehat{\pi}^{i}$ (57)
$=\varepsilon_{i}\mathop{\rm div}\nolimits^{2}\left[{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}\otimes{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right].$